OpenStudy (anonymous):

find a vector c knowing that that c is perpendicular to the vectors a=i-2j+3k , b= 2i-3j+k and c.(i+2j-7k)=100.

OpenStudy (turingtest):

let \[\vec c=x\hat i+y\hat j+z\hat k\]from last time hopefully you remember that the dot product of two perpendicular vectors is zero, so we have that\[\vec a\cdot\vec c=0\]\[\vec b\cdot\vec c=0\]and\[\vec c\cdot(\hat i+2\hat j-7\hat k)=100\]3 equations with 3 unknowns you will get a system that can be solved for x, y, and z

OpenStudy (anonymous):

yup, did it like that :P I'm only checking if i did it right :)

OpenStudy (turingtest):

well it looks like you did then :)